# Large Language Models for Design Structure Matrix Optimization

### v1

@@ Abstract
In complex engineering systems, the interdependencies among components or development activities are often modeled and analyzed using Design Structure Matrix (DSM). Reorganizing elements within a DSM to minimize feedback loops and enhance modularity or process efficiency constitutes a challenging combinatorial optimization (CO) problem in engineering design and operations. As problem sizes increase and dependency networks become more intricate, traditional optimization methods that solely use mathematical heuristics often fail to capture the contextual nuances and struggle to deliver effective solutions. In this study, we explore the potential of Large Language Models (LLMs) for helping solve such CO problems by leveraging their capabilities for advanced reasoning and contextual understanding. We propose a novel LLM-based framework that integrates network topology with contextual domain knowledge for iterative optimization of DSM element sequencing - a common CO problem. Experiments on various DSM cases show that our method consistently achieves faster convergence and superior solution quality compared to both stochastic and deterministic baselines. Notably, we find that incorporating contextual domain knowledge significantly enhances optimization performance regardless of the chosen LLM backbone. These findings highlight the potential of LLMs to solve complex engineering CO problems by combining semantic and mathematical reasoning. This approach paves the way towards a new paradigm in LLM-based engineering design optimization.
@@ 1 Introduction
Design Structure Matrix (DSM) is a well-established modeling tool in engineering design [1, 2]. It encodes relationships among elements, such as activities, components, or parameters, in a compact matrix format [2]. Reordering the sequence of nodes in a DSM to minimize feedback loops and improve modularity can improve system performance, reduce development risk, and increase design efficiency [1, 3]. Figure 1 illustrates an example of a design activity DSM before and after sequencing [4]. DSM sequencing is a typical combinatorial optimization (CO) problem, and also an NP-hard problem. As matrix sizes and dependency intricacy increase, traditional optimization methods that rely on heuristics and rule-based algorithms often struggle to find effective solutions [5, 6].
@@
    We show that incorporating contextual domain knowledge significantly improves both solution quality and convergence speed in our method. In addition, to the best of our knowledge, this is the first study to integrate LLMs and DSM for engineering optimization.
@@ 2.1 Design Structure Matrix for Engineering Design
The DSM is an essential tool that models the interactions and dependencies among elements within complex engineering systems [1]. Originating from Donald Steward’s pioneering work [2], DSMs provide a systematic framework to represent and manage dependencies among elements in complex engineering systems, using a compact matrix format. Each row and column in the matrix correspond to a specific component, activity, or parameter, while the matrix entries illustrate the state and strength of their interdependencies. DSM methodologies have evolved significantly, with researchers proposing advanced methods for managing complex engineering design processes, improving modularity, and enhancing overall system efficiency [5]. Since then, DSMs have seen significant development in both academia and industry. DSMs have been successfully applied across many disciplines for engineering design, such as automotive [18, 19], semiconductor [20], aerospace [21], factory equipment [22], and more [23, 24, 25, 26, 27].
@@
DSM can be categorized primarily into four types: component-based, team-based, activity-based, and parameter-based [28]. The first two types are static DSMs, whereas the other two are time-based DSMs. Each category addresses specific aspects of engineering design, ranging from structuring task flows and team interactions to the relationships between technical parameters of system components. Static DSMs are typically analyzed using clustering algorithms, which identify modular groupings to minimize interdependencies and facilitate parallel development. Time-based DSMs are usually analyzed using sequencing algorithms to streamline processes, reduce iterative feedback loops, and optimize the logical progression of activities.
@@
The optimization of DSMs in engineering design are typical CO problems, requiring efficient and effective methods to obtain high-quality solutions. Researchers have explored various algorithms for DSM applications, including heuristic, exact and hybrid approaches [6, 29, 30, 31, 32, 33]. These previous works have continuously advanced mathematical methods to enhance the performance and scalability of DSM optimization. However, while DSMs abstractly represent real-world engineering problems, they often encode rich contextual domain knowledge that can be valuable for guiding optimization. With the capabilities of large-scale foundational AI models, such information can be effectively incorporated into the optimization process. Thus, this paper contributes to this field by leveraging LLMs for DSM optimization, considering both semantic and mathematical reasoning.
2.2 Large Language Models for Combinatorial Optimization
@@
Combinatorial optimization problems are frequently encountered across fields, where finding an optimal solution from a finite set often drives improvements in efficiency, cost, and performance [34]. For instance, applications such as DNA barcoding and DNA assembly in synthetic biology [35], as well as job scheduling in manufacturing [36], rely heavily on effective solutions of CO problems. However, due to their NP-hard nature, these problems present substantial challenges, especially as complexity increases with larger problem sizes and more intricate dependencies. Traditionally, real-world CO problems are usually approached through the following process: the problem is first modeled formally and mathematically, then solved using specific optimization algorithms or heuristics, and finally interpreted within the context of practical engineering [37]. This separation between problem-solving and interpretation stages limits the ability to capture the contextual nuances that are often essential for optimizing real-world problems.
@@
Recent advancements have significantly enhanced the capabilities of LLMs [7, 8]. LLMs acquire extensive common-sense and domain-specific knowledge during their pretraining on large-scale datasets, including publicly available code repositories, natural language corpora, and other multimodal data. This broad foundation is further refined through fine-tuning on various downstream tasks, enabling LLMs to adapt to specific applications. State-of-the-art models such as DeepSeek, LLaMA, Claude, and GPT-series models have demonstrated remarkable performance in understanding natural language, complex reasoning, and generalizing across a wide range of tasks [38, 39, 40, 41].
@@ 3 Methodology
The process begins with the initialization of a solution that is randomly sampled from the total solution space. Each solution is then evaluated using a predefined evaluator that quantifies its quality based on specific criteria. Guided by this evaluation, the approach iteratively updates the solution base by leveraging the in-context learning capabilities of LLMs. Prompts are crafted using both the network’s mathematical representation and contextual domain knowledge in natural language descriptions, along with meta-instructions and selected historical solutions. In each iteration, the LLM generates a new candidate solution, which is then validated and evaluated. The solution, along with its evaluation results, is appended to the solution base. Once the termination condition is met, the approach outputs the best-performing solution as the final result. An overview of the proposed approach is shown in Figure 2.
@@ 3.1 Initialization and Solution Sampling
The solution base is a core component of the approach, serving three primary purposes: (1) storing all previously explored solutions along with their evaluation results, (2) providing historical solutions for few-shot learning to guide the LLM’s generation in subsequent iterations, and (3) returning the top-performing solution as the final output at the end of the optimization process.
The initialization phase begins by randomly sampling a valid solution from the entire solution space. In the context of DSM sequencing, each solution corresponds to a complete and non-repetitive permutation of nodes (see Figure 2). This initial solution is evaluated using the predefined evaluator and added to the solution base. During subsequent iterations, a sampling strategy is employed to construct a candidate solution set. Specifically, we select the top $K_{p}$ highest-scoring solutions and randomly sample additional $K_{q}$ solutions from the remaining $K_{n}-K_{p}$ entries in the solution base, where $K_{n}$ denotes the total number of stored solutions. The combined set of $K_{p}+K_{q}$ solutions is then used to construct prompts that guide the next round of LLM-based generation. The values of $K_{p}$ and $K_{q}$ are tunable hyperparameters that balance exploitation and exploration in the search process.
@@ 3.2 LLM-driven Generative Optimization
In each iteration of the optimization process, the backend LLM is prompted with a structured input that incorporates both mathematical representations of the DSM and contextual knowledge related to the given DSM. The prompt includes the following four key components:
@@
Upon receiving the complete input, the LLM generates a new candidate solution, which is then validated by a checker to ensure that it includes all nodes exactly once and adheres to the required sequencing constraints. Valid solutions are then evaluated using the predefined scoring function and subsequently appended to the solution base for use in future iterations. Details of the input information are provided in Appendix 1.
@@ 3.3 Evaluation of Generated Solutions
To evaluate solution quality, each new sequence generated by the LLM is assessed upon creation. In DSM sequencing, the objective is to reorder the rows and columns of the DSM to minimize the number of feedback loops. The resulting score is then used to rank solutions in the solution base and to guide sampling and prompt construction in subsequent iterations. For optimization problems other than DSM sequencing, the evaluation criteria can be adapted based on the structural characteristics and design objectives of the specific problem. For instance, in modularity optimization (i.e., clustering), the metric may focus on maximizing intra-module cohesion while minimizing inter-module dependencies.
When the predefined number of iterations is reached or the evaluation score meets a known optimal threshold, the iteration process is terminated. At that point, the best-performing solution identified is returned as the final output.
@@ 4.1 Data
We curated four DSM cases for experiments, which are categorized into two types following established classifications [28]: (1) Activity-based DSMs, which represent the input-output relationships among tasks or activities within a project; and (2) Parameter-based DSMs, which illustrate the precedence and interdependencies among design parameters.
@@ 4.2 Evaluation Metrics
To evaluate the quality and effectiveness of solutions generated during the optimization process, we employ the following metrics aligned with the goals of DSM sequencing tasks. The choice of metrics is consistent with prior studies [1]. Throughout the iterative optimization process, we use the number of feedback loops (i.e., non-zero entries above the diagonal in the adjacency matrix) as the primary metric to guide solution selection and LLM-driven generation.
Given a sequence, we first construct an $n\times n$ asymmetric adjacency matrix $A$ based on the edge list of the network, where $n$ denotes the number of nodes. Each binary entry $a_{ij}$ in the matrix indicates whether node $i$ depends on node $j$, with $a_{ij}=1$ representing a directed edge from $j$ to $i$. The sequencing task involves reordering the nodes such that the number of non-zero entries above the main diagonal in $A$ is minimized. These entries correspond to feedback dependencies that disrupt ideal information or task flow. Formally, given a sequence $s$, the evaluation score is computed as:
Score(s) = ∑_i=1^n ∑_j=i+1^n a_ij
@@
During performance benchmarking, we complement this primary metric with statistical reporting of mean and standard deviation over multiple trials. For convergence analysis, we record the number of unique sequences explored before reaching the optimal or near-optimal score, thereby capturing both efficiency and solution quality.
@@ 4.4 Baseline Methods
Stochastic Methods. We adopt the classic Genetic Algorithm (GA), a widely used probabilistic optimization method [48], implemented using the DEAP library [49]. Three configurations were tested to explore different optimization strategies:
@@
Detailed settings are provided in Appendix 2.
@@
Ablation Setting. To assess the role of contextual domain knowledge, we implement a variant of our method that removes all semantic information from the input prompt. In this version, the LLM is guided solely by the DSM’s topological structure, without node names or network descriptions. This setup enables us to isolate the impact of domain knowledge on optimization performance.
@@ 5.1 Performance Comparison with Benchmarks
Using the number of feedback loops as the metric (described in Section 4.2), we compared the performance of our proposed LLM-based method against both stochastic and deterministic baselines across four DSM cases, as shown in Table 2. The results demonstrate the superiority of our approach in terms of solution quality, particularly when contextual domain knowledge is incorporated.
Our methods were evaluated in three configurations: single-trial, 5-trial, and 20-trial runs, each tested with and without the inclusion of contextual domain knowledge. Among deterministic baselines, which are inherently single-trial and produce consistent outputs, our single-trial method with knowledge consistently outperformed all competitors, with the sole exception (Visibility-based method), which matched our performance in the Automobile Brake System case. In that same case, our method achieved identical best results when run under the 5-trial and 20-trial configurations.
In comparison with stochastic algorithms, our 20-trial LLM-based method consistently outperformed all GA variants across all DSM cases. It achieved the optimal solution in three out of four cases. In the remaining case (Heat Exchanger DSM), our method produced a solution that closely approached the optimum. Notably, these results were obtained within a relatively limited iteration budget of only 20 trials. Even in the 5-trial setting, our method outperformed all three GA configurations in three cases. These results validate the effectiveness of the proposed approach in producing high-quality solutions with relatively few attempts.
@@
Stochastic Methods [^1] Consistent with Figure 3, all three GA settings show optimization performance with number of unique sequences explored = 10,000.
GA (Exploration-focused setting) 6.0±0.0 8.1±0.3 5.7±1.0 3.8±0.7
GA (Exploitation-focused setting) 7.4±1.5 9.9±1.4 7.6±1.4 6.6±1.7
GA (Balanced setting) 6.0±0.0 8.4±0.5 6.2±1.2 4.2±1.2

### v2

# Large Language Models for Combinatorial Optimization of Design Structure Matrix

@@ Abstract
In complex engineering systems, the dependencies among components or development activities are often modeled and analyzed using Design Structure Matrix (DSM). Reorganizing elements within a DSM to minimize feedback loops and enhance modularity or process efficiency constitutes a challenging combinatorial optimization (CO) problem in engineering design and operations. As problem sizes increase and dependency networks become more intricate, traditional optimization methods that rely solely on mathematical heuristics often fail to capture the contextual nuances and struggle to deliver effective solutions. In this study, we explore the potential of Large Language Models (LLMs) to address such CO problems by leveraging their capabilities for advanced reasoning and contextual understanding. We propose a novel LLM-based framework that integrates network topology with contextual domain knowledge for iterative optimization of DSM sequencing - a common CO problem. Experiments on various DSM cases demonstrate that our method consistently achieves faster convergence and superior solution quality compared to both stochastic and deterministic baselines. Notably, incorporating contextual domain knowledge significantly enhances optimization performance regardless of the chosen LLM backbone. These findings highlight the potential of LLMs to solve complex engineering CO problems by combining semantic and mathematical reasoning. This approach paves the way towards a new paradigm in LLM-based engineering design optimization.
@@ 1 Introduction
Design Structure Matrix (DSM) is a well-established modeling tool in engineering design [1, 2]. It encodes relationships among elements, such as activities, components, or parameters, in a compact matrix format [2]. Reordering the sequence of nodes in a DSM to minimize feedback loops and improve modularity can improve system performance, reduce development risk, and increase design efficiency [3, 1]. Consider a simple DSM with three tasks: A, B, and C. If A must precede B, and B must precede C, then a valid sequence [A, B, C] has no feedback loops. However, if the sequence is [B, A, C], then the dependency from A to B becomes a “feedback loop,” reducing efficiency.
Figure 1 illustrates a real-world example of a design activity DSM before and after sequencing [4]. DSM sequencing is a typical combinatorial optimization (CO) problem, and also an NP-hard problem. As matrix sizes and dependency intricacy increase, traditional optimization methods that rely on heuristics and rule-based algorithms often struggle to find effective solutions [5, 6].
@@
    We show that incorporating domain knowledge significantly improves both solution quality and convergence speed in our method. To the best of our knowledge, this is the first study to integrate LLMs and DSM for engineering optimization.
@@ 2.1 Design Structure Matrix for Engineering Design
The DSM is an essential tool that models the interactions and dependencies among elements within complex engineering systems [1]. Originating from Steward’s pioneering work [2], DSMs provide a systematic framework to represent and manage dependencies among elements in complex engineering systems, using a compact matrix format. Each row and column in the matrix correspond to a specific component, activity, or parameter, while the matrix entries illustrate the state and strength of their dependencies. DSM methodologies have evolved significantly, with researchers proposing advanced methods for managing complex engineering design processes, improving modularity, and enhancing overall system efficiency [5]. Since then, DSMs have seen significant development in both academia and industry. DSMs have been successfully applied across many disciplines for engineering design, such as automotive [18, 19], semiconductor [20], aerospace [21], factory equipment [22], manufacturing systems [23, 24], and more [25, 26, 27, 28, 29].
@@
DSM can be categorized primarily into four types: component-based, team-based, activity-based, and parameter-based [30]. The first two types are static DSMs, whereas the other two are time-based DSMs. Each category addresses specific aspects of engineering design, ranging from structuring task flows and team interactions to the relationships between technical parameters of system components. Static DSMs are typically analyzed using clustering algorithms, which identify modular groupings to minimize dependencies and facilitate parallel development. Time-based DSMs are usually analyzed using sequencing algorithms to streamline processes, reduce iterative feedback loops, and optimize the logical progression of activities.
@@
The optimization of DSMs in engineering design are typical CO problems, requiring efficient and effective methods to obtain high-quality solutions. Researchers have explored various algorithms for DSM applications, including heuristic, exact and hybrid approaches [31, 32, 33, 34, 6, 35]. For example, Meier et al. [36] proposed the Ordering messy Genetic Algorithm (OmeGA), which integrates modified GA and local search to enhance DSM sequencing. Qian et al. [6] presented a hybrid heuristic-exact method based on block folding, which shows progressive improvement and performs efficiently on the DSM sequencing problem. These previous works have continuously advanced mathematical methods to enhance the performance and scalability of DSM optimization. However, while DSMs abstractly represent real-world engineering problems, they often encode rich contextual domain knowledge that can be valuable for guiding optimization. With the capabilities of large-scale foundation models, such information can be effectively incorporated into the optimization process. This paper contributes to this field by leveraging LLMs for DSM optimization, considering both semantic and mathematical reasoning.
2.2 Large Language Models for Combinatorial Optimization (CO)
@@
Combinatorial optimization problems are frequently encountered across diverse engineering domains, where finding an optimal solution from a finite set often drives improvements in efficiency, cost, and performance [37]. For instance, applications such as DNA barcoding and DNA assembly in synthetic biology [38], as well as job scheduling in manufacturing [39], rely heavily on effective solutions of CO problems. However, due to their NP-hard nature, these problems present substantial challenges, especially as complexity increases with larger problem sizes and more intricate dependencies. Traditionally, real-world CO problems are usually approached through the following process: the problem is first modeled formally and mathematically, then solved using specific optimization algorithms or heuristics, and finally interpreted within the context of practical engineering [40]. This separation between problem-solving and interpretation stages limits the ability to capture the contextual nuances that are often essential for optimizing real-world problems.
@@
Recent advancements have significantly enhanced the capabilities of LLMs [7, 8]. LLMs acquire extensive common-sense and domain-specific knowledge during their pretraining on large-scale datasets, including publicly available code repositories, natural language corpora, and other multimodal data. This broad foundation is further refined through fine-tuning on various downstream tasks, enabling LLMs to adapt to specific applications. State-of-the-art LLMs such as DeepSeek, LLaMA, Claude, and GPT-series models have demonstrated remarkable performance in understanding natural language, complex reasoning, and generalizing across a wide range of tasks [41, 42, 43, 44].
@@ 3 Methodology
The process begins with the initialization of a valid solution that is randomly sampled from the total solution space. Each solution is then evaluated using a predefined evaluator that quantifies its quality based on specific criteria. Guided by this evaluation, the approach iteratively updates the solution base by leveraging the in-context learning capabilities of LLMs. Prompts are crafted using both the network’s mathematical representation and contextual domain knowledge in natural language descriptions, along with meta-instructions and selected historical solutions. In practice, the DSM data used in this process are constructed from documented system architectures and task dependencies, which may be extracted from prior studies, design specifications, or expert elicitation. In each iteration, the LLM generates a new candidate solution, which is then validated and evaluated. The solution, along with its evaluation results, is appended to the solution base. Once the termination condition is met, the approach outputs the best-performing solution as the final result. An overview of the proposed approach is shown in Figure 2.
@@ 3.1 Initialization and Solution Sampling
The initialization begins by randomly sampling a valid solution from the entire solution space. In the context of DSM sequencing, each solution corresponds to a complete and non-repetitive permutation of nodes. For instance, in a DSM with nodes A, B, C, one possible candidate sequence is [B, A, C]. This initial solution is evaluated using the predefined evaluator and added to the solution base. The solution base is a key component of the approach, serving three primary purposes: (1) storing all previously explored solutions along with their evaluation results ([Solution, Score] pairs), (2) providing historical solutions for few-shot learning to guide the LLM’s generation in subsequent iterations, and (3) returning the top-performing solution as the final output at the end of the optimization process.
During subsequent iterations, a sampling strategy is employed to construct a candidate solution set for LLM in-context learning. Specifically, we select the top $K_{p}$ highest-scoring solutions and randomly sample additional $K_{q}$ solutions from the remaining $K_{n}-K_{p}$ entries in the solution base, where $K_{n}$ denotes the total number of stored solutions. The combined set of $K_{p}+K_{q}$ solutions is then used to construct prompts that guide the next round of LLM-based generation. The values of $K_{p}$ and $K_{q}$ are tunable hyperparameters that balance exploitation and exploration in the search process. These historical solutions are formatted as a list of [Solution, Score] pairs in the prompt.
Figure 3: Structure of the input to LLMs for DSM sequencing optimization
@@ 3.2 LLM-driven Generative Optimization
In each iteration of the optimization process, the backend LLM is prompted with a structured input information that incorporates both mathematical representations of the DSM and contextual knowledge related to the given DSM. The prompt includes the following four key components:
@@
Upon receiving the complete input, the LLM generates a new candidate solution, which is then validated by a checker to ensure that it includes all nodes exactly once and adheres to the required sequencing constraints. Valid solutions are then evaluated using the predefined scoring function and subsequently appended to the solution base for use in future iterations.
Figure 3 illustrates an example prompt, where each component is explicitly represented: meta-instructions define the task and output requirements; contextual domain knowledge specifies node descriptions and system context; typological information provides the DSM’s node and edge list; and selected historical solutions offer prior orderings with scores. More details refer to Appendix 1.
@@ 3.3 Evaluation of Generated Solutions
To evaluate solution quality, each new sequence generated by the LLM is assessed upon creation. In general, a wide range of objectives have been proposed for DSM sequencing in prior research, depending on whether the emphasis is on reducing iteration, improving concurrency, lowering cost, or enhancing modularity. Examples include minimizing feedbacks or weighted feedbacks [48, 2], minimizing feedback length [49], and more [50, 51]. In this study, we adopt the most widely used objective: minimizing the number of feedback loops (i.e., non-zero entries above the diagonal of the DSM adjacency matrix). This criterion has been consistently employed in the literature as a baseline indicator of process rework, making it appropriate for evaluating the feasibility of an LLM-driven optimization approach.
The resulting score is then used to rank solutions in the solution base and to guide sampling and prompt construction in subsequent iterations. For optimization problems other than DSM sequencing, the evaluation criteria can be adapted based on the structural characteristics and design objectives of the specific problem. For instance, in modularity optimization (i.e., partitioning task), the metric may focus on maximizing intra-module cohesion while minimizing inter-module dependencies. When the predefined number of iterations is reached or the evaluation score meets a known optimal threshold, the iteration process is terminated. The best-performing solution is then returned as the final output.
@@ 4.1 Data
We collected four DSM cases for experiments, which are categorized into two types following established classifications [30]: (1) Activity-based DSMs, which represent the input-output relationships among tasks or activities within a project; and (2) Parameter-based DSMs, which illustrate the precedence and dependencies among design parameters.
@@ 4.2 Evaluation Metrics
DSM sequencing can be assessed under multiple objectives depending on the design context. Prior studies have considered criteria such as reducing iteration, enhancing concurrency, minimizing development time and cost, or improving modularity of the system [51, 49, 48, 2]. In this study, we adopt the minimization of feedback loops as the evaluation metric, as it is the widely adopted measure in the DSM literature [2]. It is noteworthy that the sequence with the fewest feedbacks is not always the practically optimal one, but it can serve as a reasonable foundation for evaluating our approach.
To evaluate a given sequence, we first construct an $n\times n$ asymmetric adjacency matrix $A$ based on the edge list of the network, where $n$ denotes the number of nodes. Each binary entry $a_{ij}$ in the matrix indicates whether node $i$ depends on node $j$, with $a_{ij}=1$ representing a directed edge from $j$ to $i$. The sequencing task involves reordering the nodes such that the number of non-zero entries above the main diagonal in $A$ is minimized. These entries correspond to feedback dependencies that disrupt ideal information or task flow. Formally, given a sequence $s$, the evaluation score is computed as:
$\text{Score}(s)=\sum_{i=1}^{n}\sum_{j=i+1}^{n}a_{ij}$
@@
During performance benchmarking, we complement this primary metric with statistical reporting of mean and standard deviation over multiple trials. For convergence analysis, we record the number of unique sequences explored before reaching the optimal score, thereby capturing both efficiency and solution quality.
@@ 4.4 Baseline Methods
Stochastic Methods. We adopt the classic Genetic Algorithm (GA), a widely used probabilistic optimization method [55]. Three configurations were tested to explore different optimization strategies:
@@
In addition, we include the Ordered Messy Genetic Algorithm (OmeGA), which extends the standard GA with purpose-specific enhancements for DSM sequencing [36]. OmeGA first decomposes the DSM into strongly connected components (SCCs), applies topological sorting for inter-block ordering, and then performs genetic optimization with integrated local search within each block. Detailed settings are described in Appendix 2.
@@
Ablation Setting. To assess the role of contextual domain knowledge, we implement a variant of our method that removes all semantic information from the input prompt (the “without knowledge” setting in Table 2, Table 3, and Figure 4). In this version, the LLM is guided solely by the DSM’s topological information, without node names or network descriptions. This setup enables us to isolate the impact of domain knowledge on optimization performance.
@@ 5.1 Performance Comparison with Benchmarks
Using the number of feedback loops as the metric (described in Section 4.2), we compared the performance of our proposed LLM-based method against both stochastic and deterministic baselines across four DSM cases, as shown in Table 2. The entries represent the number of feedback loops. Note that fractional values arise because the results are averaged over multiple independent runs, and the ± terms indicate one standard deviation, reflecting variability in performance.
Our methods were evaluated in three configurations: single-trial, 5-trial, and 20-trial runs, each tested with and without the inclusion of contextual domain knowledge. Among deterministic baselines, which are inherently single-trial and produce consistent outputs, our single-trial method with knowledge consistently outperformed all baseline methods, with the sole exception of the Visibility-based method, which matched our performance in the Automobile Brake System DSM case. In that same case, our method achieved identical best results when run under the 5-trial and 20-trial configurations. When compared to stochastic baselines, the 20-trial LLM-based method matched or outperformed all GA variants and OmeGA in three of the four cases. Although OmeGA reached the optimal solution in the Heat Exchanger DSM, our method delivered strong performance surpassing all GA variants despite being limited to only 20 trials. In the remaining three DSM cases, our LLM-based method consistently found the optimal solutions within 20 trials, demonstrating its strong effectiveness under constrained iteration budgets. With only five trials, our method delivered competitive results compared to all stochastic methods in most cases. These findings highlight the efficiency of the proposed approach in generating high-quality solutions with minimal search effort.
@@
Stochastic Methods [^1] For stochastic methods, GA-500 denotes the performance after 500 solutions have been explored (evaluated), and GA-10K denotes the performance after 10,000 solutions, consistent with Figure 4. Other GA variants are defined similarly.
GA-500 (Exploration-focused) 7.9±1.1 9.1±0.9 7.2±0.9 5.2±0.9
GA-10K (Exploration-focused) 6.0±0.0 8.1±0.3 5.7±1.0 3.8±0.7
GA-500 (Exploitation-focused) 14.1±2.9 16.7±2.1 12.5±2.2 11.4±2.5
GA-10K (Exploitation-focused) 7.4±1.5 9.9±1.4 7.6±1.4 6.6±1.7
GA-500 (Balanced) 8.8±1.3 10.9±1.0 8.9±1.7 6.8±1.6
GA-10K (Balanced) 6.0±0.0 8.4±0.5 6.2±1.2 4.2±1.2
OmeGA-500 [36] 6.3±0.9 8.0±0.0 3.5±0.5 3.0±0.0
OmeGA-10K [36] 6.0±0.0 8.0±0.0 3.0±0.0 3.0±0.0